# UniTable: Towards a Unified Framework for Table Structure Recognition via Self-Supervised Pretraining

### v1

@@ 4 Experiments
4.1 Implementation Details
Architecture. We have trained two model variants: (1) a base model with 30M parameters including 4 encoder layers, 8 attention heads with a hidden size of 512, (2) a large model with 125M parameters including 12 encoder layers, 12 attention heads with a hidden size of 768. Both base and large models have a task decoder of 4 decoder layers. The maximum token sequence length is 512 for table structure, 1024 for cell bbox, and 200 for cell content, as we find such settings satisfy most of the tables.
Training and inference We have pretrained the VQ-VAE on 1M and 2M tabular images. The 1M VQ-VAE is trained on PubTabNet (Zhong et al., 2020) and SynthTabNet (Nassar et al., 2022), and the extra 1M datasets for training 2M VQ-VAE are PubTables-1M (Smock et al., 2022) and TableBank (Li et al., 2020). In Sec. 4.3, we present the finetuning results of 2M VQ-VAE for comparing with SOTA methods. In Sec. 5.1, we ablate the effectiveness and scalability of SSP on both 1M and 2M VQ-VAE. All models are trained with the AdamW optimizer (Loshchilov & Hutter, 2017). We employ a cosine learning rate scheduler with a linear warmup. All models are trained for 24 epochs for a fair comparison. We apply teacher forcing during training and employ greedy decoding at inference time.
@@ 4.2 Evaluation Metrics
car was first proposed in ICDAR2013 competition (Göbel et al., 2013) and improved by ICDAR2019 competition (Gao et al., 2019). It aligns the the predicted bbox with the groundtruth bbox for each table cell based on intersection over union (IoU) and generates a list of adjacency relations between a non-empty cell and its nearest horizontal and vertical neighbors. The precision, recall, and F1 score are computed based on this converted 1-D adjacency relation list.
COCO average prevision (AP) (Lin et al., 2014) is a widely used metric for generic object detection, which has been reported in other work for evaluating table cell detection. We use the COCO evaluation toolkit [^1] https://github.com/cocodataset/cocoapi and report mean (mAP), $\text{AP}_{50}$, and $\text{AP}_{75}$.
@@
4.3 Datasets and Results
We evaluate UniTable on five of the largest publicly available TSR datasets as shown in Table 1. Comparing UniTable with the prior SOTA on each dataset across all available tasks, we achieve new SOTA on four out of the five datasets even without training with task-specific loss (Huang et al., 2023) or relying on external PDF for text extraction and bbox post-processing (Nassar et al., 2022). Below is an introduction of each dataset and comparisons with previous methods.
ICDAR 2019 B2 Modern (IC19B2M) (Gao et al., 2019) was originated from ICDAR 2019 table competition. The dataset has two subsets for TSR, archival and modern, and only the modern subset has table cell content bbox annotations. The competition computes CAR F1 score at IoU$\in[0.6,0.7,0.8,0.9]$ and ranks method by weighted average F1 (WAvg. F1):
@@ 4.3 Datasets and Results
We evalute on all 100 test tables from the modern subset and report both WAvg. F1 and IoU 0.6 as in other work. Our method significantly improves the previous SOTA GTE (Zheng et al., 2021) by a large margin.
@@
PubTabNet (Zhong et al., 2020) contains 509k images of heterogeneous tables extracted from the medical scientific articles. It is the first large-scale TSR dataset that provides annotations (in HTML format) of table cell bbox, table structure, and table cell content. PubTabNet measures the table cell bbox by COCO $\text{AP}_{50}$, table structure by S-TEDS, and full table including both structure and cell content by TEDS. The authors of PubTabNet also developed the EDD model, which consisted of a CNN encoder and dual recurrent neural network (RNN) decoders for predicting table structure and cell content, respectively. TableFormer (Nassar et al., 2022) improved EDD by replacing the cell content decoder with a cell bbox decoder and extracted all contents from the PDF corresponding to the tabular image. VAST (Huang et al., 2023) added an auxiliary visual-alignment loss while training the cell bbox decoder and achieved previous SOTA on all three metrics. Our unified training framework with SSP achieves the new SOTA on all tasks even without leveraging any external PDF as PDF corresponding to the tabular image may not always exist. Specifically, both UniTable-base and UniTable-large outperform VAST on $\text{AP}_{50}$ by more than 3 percentage points (pp), which confirms the effectiveness of converting the cell bbox detection to language modeling.
@@
FinTabNet (Zheng et al., 2021) is a dataset containing 113k tables from the annual reports of the S&P 500 companies in PDF format. The major challenge of this dataset is that financial tables largely differ from scientific and government document tables in that the former has fewer graphical lines, larger gaps within each table, and more color variations. Thus, FinTabNet mainly evaluates table structure prediction accuracy, i.e., S-TEDS. VAST trained the model with an auxiliary supervised signal and achieved the previous SOTA. We achieve the new SOTA by leveraging SSP and finetuning without a task-specific loss objective. The performance gain from base to large shows that our method scales with model parameters.
@@
SynthTabNet (Nassar et al., 2022) is a large-scale synthetically generated dataset that offers control over 1) dataset size, 2) table structure, 3) table style, and 4) content type. The dataset aims to overcome the limitations of PubTabNet and FinTabNet, which are skewed table distributions towards simpler tables, limited variance in appearance styles, and restricted cell content domains. Thus, SynthTabNet is organized into 4 subsets of 150k tables (600k in total), namely Finance, PubTabNet, Marketing, and Sparse. The first two mimic the appearance of FinTabNet and PubTabNet but encompass more complex table structures. Marketing adopts a colorful appearance with high contrast that resembles real-world marketing tables, and Sparse contains tables with sparse content. The authors of SynthTabNet propose TableFormer as a baseline. TableFormer predicts table structure and cell bbox and relies on external PDF to extract cell content, so $\text{AP}_{50}$ and S-TEDS are the evaluation metrics. Both our base and large models outperform previous SOTA on $\text{AP}_{50}$ and S-TEDS. Since SynthTabNet is a comprehensive dataset that can thoroughly evaluate the model under different table configurations, we use it for ablations and present results on four subsets separately in Sec. 5.
@@
PubTables-1M (Smock et al., 2022) aims to overcome the groundtruth inconsistency observed in prior datasets using a new canonicalization procedure. The dataset has 947k tables annotated with bbox and text within each bbox. The dataset differs from previous datasets that the bbox is word-wise instead of cell-wise, thus each cell can have more than one bbox. Besides, since all annotations are in bbox format and no table structure labels, e.g., HTML, are provided, the baseline DETR trained by the dataset creators report their performance in detection metrics, e.g., $\text{AP}_{50}$ and $\text{AP}_{75}$. UniTable also achieves competitive results compared with the baseline DETR. After visualizing our predicted bbox, we find our model normally predicts more bboxes (longer sequence) than the groundtruth. We delve deep into the dataset annotations and discover previously unacknowledged inconsistencies in table annotations. We conjecture the SOTA performance of DETR model on PubTables-1M may be due to overfitting to the training set.
@@ 4.4 Using UniTable in Practice
We have also finetuned our UniTable-large across multiple datasets and released a Jupyter Notebook as a demo of our inference pipeline. A user can simply pass a table screenshot through our notebook and obtain a fully digitalized HTML table. The table structure is trained across PubTabNet, FinTabNet, and SynthTabNet, the cell bbox is trained across PubTabNet and SynthTabNet, and the cell content is trained across PubTabNet, SynthTabNet, and PubTables-1M. Appendix A visualizes several table examples in practice.
@@ 5.1 Effectiveness and Scalability of SSP
Table 2: Effectiveness and scalability of SSP on all four subsets of SynthTabNet. Without SSP, the model performance suffers, and increasing the model complexity from base to large barely improves the performance. Pretraining the visual encoder on 1M tabular images provides an average increase of 14.40 pp. Pretraining on 2M images continues to increase the performance by 0.74 pp. Here we present S-TEDS of the table structure prediction, and the same trend also applies to other tasks as elaborated in Appendix B.
@@
We ablate the effectiveness and scalability of SSP on all four subsets of SynthTabNet since SynthTabNet is a large-scale dataset that can comprehensively evaluate the model under different table configurations. Table 2 presents the table structure task, and the same trend also applies to other tasks as elaborated in Appendix B. Comparing row “No SSP” and “SSP 1M” or “SSP 2M”, both the base and the large models have benefited significantly from the SSP. Specifically, Marketing and Sparse are two of the most challenging subsets. Marketing has a colorful background with high-contrast texts, and Sparse contains tables with sparse content. These variations make it challenging to accurately predict the HTML table structure tags and bbox surrounding the cell content. Without SSP, the model performance suffers, and increasing the model complexity from base to large barely improves the performance. Instead, after pretraining the visual encoder on 1M tabular images, both the base and the large models have an average increase of 14.40 pp. Note the performance also scales along with the pretraining dataset size. When the pretraining dataset increases from 1M to 2M, the average performance continues to increase by 0.74 pp. Finally, all large models have an average performance gain of 1.51 pp over the base models.
@@ 5.2 Generalization of the Unified Training Objective
Table 3: UniTable’s unified training objective applies to both linear projection Transformer and hybrid CNN-Transformer architectures conventionally used in TSR. Results on all four subsets of the SynthTabNet for table structure prediction evaluated with S-TEDS, and the same conclusion also applies to other tasks as elaborated in Appendix C.
@@
We showcase that our unified training objective, language modeling, not only works on models with linear projection Transformer, but also works with the hybrid CNN-Transformer architecture used in the TSR literature. Table 3 presents the results on all four subsets of the SynthTabNet similar to the settings in Sec. 5.1. The main difference is that the linear projection in the visual encoder is replaced by a ResNet-18 CNN backbone. Such a modification leads to an increase of 12M parameters. The performance of this hybrid CNN-Transformer architecture is roughly on par with the linear projection model initialized from the SSP 1M tabular images, which shows that our training objective is agnostic to the choice of architecture. The significant performance gap between hybrid CNN-Transformer and linear projection Transformer trained from scratch verifies the observation identified in the previous literature (Peng et al., 2023). Though hybrid CNN-Transformer has also achieved competitive results, we still recommend using the linear projection Transformer because of 1) capability of leveraging the power of SSP, 2) architectural compliance with VLM in natural image domain, and 3) the performance of hybrid CNN-Transformer is still worse than the SSP on 2M tabular images even with more total parameters. Though prior work has started to propose SSP for hybrid CNN-Transformer, this direction is still in its early stage as it is challenging in heavy computational cost and pretraining-finetuning discrepancy (Gao et al., 2022). Moreover, most work only demonstrates the capability in discriminative tasks rather than the generative tasks used in the TSR domain. Besides, existing SOTA VLMs, e.g., BLIP-2 (Li et al., 2023) and LLaVA (Liu et al., 2023), still employ the linear projection and leave the Transformer to learn the interactions between different patches. As we believe the compliance of architecture is an important step towards generic VLMs, our success in replacing the CNN backbone with linear projection and leveraging a task-agnostic language modeling loss is a cornerstone of incorporating TSR in modern VLM training.
@@ 5.3 Why does SSP work?
((a)) Red patches represent background.
((b)) Red patches represent separations in table.
@@
Figure 2: The key reason that SSP works is because each tokens have visual semantics and the codebook shows a a fine-grained categorization to represent the implicit conventions in the table. The codebook used in SSP has learned to represent abstract concepts by using different groups of tokens to represent different concepts: (a) empty white background and (b) separations within a table. Red highlights the token indices under investigation. Appendix D provides a zoomed-in version of these images labeled with token indices. We lay the color patches over the original tabular image to present the selected token indices from the 2M VQ-VAE.
@@
During SSP, the visual encoder is trained to fill in the masked tabular image by selecting visual tokens from the VQ-VAE codebook. Thus, we visualize the selected token indices and analyze whether the tokens have visual semantics. Fig. 2 presents the original table overlaid by the selected token indices from the 2M VQ-VAE codebook. We highlight the token indices under investigation with red patches. Appendix D provides a zoomed-in version of these images labeled with token indices. The input image size is $448\times 448$ and the patch size is $16\times 16$, so each image has $28\times 28$ tokens chosen from the 16384 entries in the codebook. First, we highlight the tokens representing the blank white background. Comparing the highlighted (red) and not-highlighted regions, we observe that the highlighted region has covered most of the blank space inside the table, and the not-highlighted region has formulated a convex hull tightly surrounding either the texts or separation lines. Next, we highlight the tokens representing the separation lines. Comparing the highlighted (red) and not-highlighted regions, we observe that the red regions align with all the separations within the table, e.g., line separations and color shading separations. Taking a closer look at the indices in Fig. 4 and 5, we find the codebook has learned to represent an abstract concept by assigning multiple tokens, where each reflects a different scenario. For example, token “111” represents the separation above the gray color shading and token “964” represents the one below, and token “15282” represents the separation above a bold horizontal line and “10807” represents the one below. A certain type of separation also has multiple choices depending on the portion of the line within the patch, e.g., token “14181”, “8714”, and “10807” all representing the “below a bold horizontal line”, but the differences lie in the line thickness and amount of black inside the image patch. Such a fine-grained categorization also explains why the codebook needs so many tokens to represent the implicit conventions created by humans in the table.
6 Related Work
Prior work on TSR is scattered and has mainly centered around complex task-specific combinations of available inputs and tools. The earliest methods were based on splitting and merging, where the first stage detected rows, columns, and text blocks, and the second stage merged text blocks to restore the structure, e.g., DeepDeSRT (Schreiber et al., 2017) and TableNet (Paliwal et al., 2019). Another line of methods first detects the cells and then classifies the row and column relationships based on graphs, e.g., FLAG-NET (Liu et al., 2021a) and Hetero-TSR (Liu et al., 2022). Recent researchers have also started to leverage image-to-text generation for table structure prediction. This method treats the table structure as a sequence and adopts an end-to-end image-to-text paradigm. Deng, et al. (Deng et al., 2019) employed a hybrid CNN-LSTM architecture to generate the LaTeX code of the table. Zhong, et al. (Zhong et al., 2020) introduced an encoder-dual-decoder (EDD) architecture in which two RNN-based decoders were responsible for logical and cell content, respectively. Both TableFormer (Nassar et al., 2022) and TableMaster (Ye et al., 2021) enhanced the EDD decoder with a transformer decoder and included a regression decoder to predict the bounding box instead of the content. VAST (Huang et al., 2023) proposed an auxiliary visual alignment loss to ensure that the logical representation of the non-empty cells contains more local visual details. Our goal differs from existing approaches as we focus on providing a training framework that unifies both the training paradigm and the training objective of TSR.

### v2

@@ 4 Experiments
4.1 Implementation
Architecture. We have trained two model variants: (1) a base model with 30M parameters including 4 encoder layers, 8 attention heads with a hidden size of 512, (2) a large model with 125M parameters including 12 encoder layers, 12 attention heads with a hidden size of 768. Both base and large models have a task decoder of 4 decoder layers. The maximum token sequence length is 512 for table structure, 1024 for cell bbox, and 200 for cell content, as we find such settings satisfy most tables.
Training and inference We have pretrained the VQ-VAE on 1M and 2M tabular images. The 1M VQ-VAE is trained on PubTabNet [42] and SynthTabNet [28], and the extra 1M datasets for training 2M VQ-VAE are PubTables-1M [36] and TableBank [18]. In Sec. 4.3, we present the finetuning results of 2M VQ-VAE for comparing with SOTA methods. In Sec. 5.1, we ablate the effectiveness and scalability of SSP on both 1M and 2M VQ-VAE. All models are trained with the AdamW optimizer [23]. We employ a cosine learning rate scheduler with a linear warmup. All models are trained for 24 epochs for a fair comparison. We apply teacher forcing during training and employ greedy decoding at inference time.
@@ 4.2 Evaluation Metrics
car was first proposed in ICDAR2013 competition [10] and improved by ICDAR2019 competition [8]. It aligns the the predicted bbox with the groundtruth bbox for each table cell based on intersection over union (IoU) and generates a list of adjacency relations between a non-empty cell and its nearest horizontal and vertical neighbors. The precision, recall, and F1 score are computed based on this converted 1-D adjacency relation list.
COCO average prevision (AP) [19] is a widely used metric for generic object detection, which has been reported in other work for evaluating table cell detection. We use the COCO evaluation toolkit [^2] https://github.com/cocodataset/cocoapi and report mean (mAP), $\text{AP}_{50}$, and $\text{AP}_{75}$.
@@
4.3 Results on Datasets
We evaluate UniTable on five of the largest publicly available TR datasets as shown in Table 1. Comparing UniTable with the prior SOTA on each dataset across all available tasks, we achieve new SOTA on four out of the five datasets even without training with task-specific loss [12] or relying on external PDF for text extraction and bbox post-processing [28]. Below is an introduction of each dataset and comparisons with previous methods.
ICDAR 2019 B2 Modern (IC19B2M) [8] was originated from ICDAR 2019 table competition. The dataset has two subsets for TR, archival and modern, and only the modern subset has table cell content bbox annotations. The competition computes CAR F1 score at IoU$\in[0.6,0.7,0.8,0.9]$ and ranks method by weighted average F1 (WAvg. F1):
@@ 4.3 Results on Datasets
We evalute on all 100 test tables from the modern subset and report both WAvg. F1 and IoU 0.6 as in other work. Our method significantly improves the previous SOTA GTE [41] by a large margin.
@@
PubTabNet [42] contains 509k images of heterogeneous tables extracted from the medical scientific articles. It is the first large-scale TR dataset that provides annotations (in HTML format) of table cell bbox, table structure, and table cell content. PubTabNet measures the table cell bbox by COCO $\text{AP}_{50}$, table structure by S-TEDS, and full table including both structure and cell content by TEDS. The authors of PubTabNet also developed the EDD model, which consisted of a CNN encoder and dual recurrent neural network (RNN) decoders for predicting table structure and cell content, respectively. TableFormer [28] improved EDD by replacing the cell content decoder with a cell bbox decoder and extracted all contents from the PDF corresponding to the tabular image. VAST [12] added an auxiliary visual-alignment loss while training the cell bbox decoder and achieved previous SOTA on all three metrics. Our unified training framework with SSP achieves the new SOTA on all tasks even without leveraging any external PDF as PDF corresponding to the tabular image may not always exist. Specifically, both UniTable-base and UniTable-large outperform VAST on $\text{AP}_{50}$ by more than 3 percentage points (pp), which confirms the effectiveness of converting the cell bbox detection to language modeling.
@@
FinTabNet [41] is a dataset containing 113k tables from the annual reports of the S&P 500 companies in PDF format. The major challenge of this dataset is that financial tables largely differ from scientific and government document tables in that the former has fewer graphical lines, larger gaps within each table, and more color variations. Thus, FinTabNet mainly evaluates table structure prediction accuracy, i.e., S-TEDS. VAST trained the model with an auxiliary supervised signal and achieved the previous SOTA. We achieve the new SOTA by leveraging SSP and finetuning without a task-specific loss objective. The performance gain from base to large shows that our method scales with model parameters.
@@
SynthTabNet [28] is a large-scale synthetically generated dataset that offers control over 1) dataset size, 2) table structure, 3) table style, and 4) content type. The dataset aims to overcome the limitations of PubTabNet and FinTabNet, which are skewed table distributions towards simpler tables, limited variance in appearance styles, and restricted cell content domains. Thus, SynthTabNet is organized into 4 subsets of 150k tables (600k in total), namely Finance, PubTabNet, Marketing, and Sparse. The first two mimic the appearance of FinTabNet and PubTabNet but encompass more complex table structures. Marketing adopts a colorful appearance with high contrast that resembles real-world marketing tables, and Sparse contains tables with sparse content. The authors of SynthTabNet propose TableFormer as a baseline. TableFormer predicts table structure and cell bbox and relies on external PDF to extract cell content, so $\text{AP}_{50}$ and S-TEDS are the evaluation metrics. Both our base and large models outperform previous SOTA on $\text{AP}_{50}$ and S-TEDS. Since SynthTabNet is a comprehensive dataset that can thoroughly evaluate the model under different table configurations, we use it for ablations and present results on four subsets separately in Sec. 5.
@@
PubTables-1M [36] aims to overcome the groundtruth inconsistency observed in prior datasets using a new canonicalization procedure. The dataset has 947k tables annotated with bbox and text within each bbox. The dataset differs from previous datasets that the bbox is word-wise instead of cell-wise, thus each cell can have more than one bbox. Besides, since all annotations are in bbox format and no table structure labels, e.g., HTML, are provided, the baseline DETR trained by the dataset creators report their performance in detection metrics, e.g., $\text{AP}_{50}$ and $\text{AP}_{75}$. UniTable also achieves competitive results compared with the baseline DETR. Visualizing the bbox predictions shows that our model predicts more bboxes (longer sequence) than the groundtruth, motivating us to delve deep into the dataset annotations and discover previously unacknowledged inconsistencies in table annotations. We conjecture the SOTA performance of DETR model on PubTables-1M may be due to overfitting to the training set. Sec. F.1 describes the details of these dataset annotation issues.
@@ 4.4 Using UniTable in Practice
We have also finetuned our UniTable-large across multiple datasets and released a Jupyter Notebook as a demo of our inference pipeline. A user can simply pass a table screenshot through our notebook and obtain a fully digitalized HTML table. The table structure is trained across PubTabNet, FinTabNet, and SynthTabNet, the cell bbox is trained across PubTabNet and SynthTabNet, and the cell content is trained across PubTabNet, SynthTabNet, and PubTables-1M. We have provided a public API hosted on HuggingFace to facilitate the access to our UniTable: https://poloclub.github.io/magic-table/. Appendix B visualizes several table examples in practice.
@@ 5.1 Effectiveness and Scalability of SSP
Table 2: Effectiveness and scalability of SSP on all four subsets of SynthTabNet. Without SSP, the model performance suffers, and increasing the model complexity from base to large barely improves the performance. Pretraining the visual encoder on 1M tabular images provides an average increase of 14.40 pp. Pretraining on 2M images continues to increase the performance by 0.74 pp. Here we present S-TEDS of the table structure prediction, and the same trend also applies to other tasks as elaborated in Appendix C.
@@
We ablate the effectiveness and scalability of SSP on all four subsets of SynthTabNet since SynthTabNet is a large-scale dataset that can comprehensively evaluate the model under different table configurations. Table 2 presents the table structure task, and the same trend also applies to other tasks as elaborated in Appendix C. Comparing row “No SSP” and “SSP 1M” or “SSP 2M”, both the base and the large models have benefited significantly from the SSP. Specifically, Marketing and Sparse are two of the most challenging subsets. Marketing has a colorful background with high-contrast texts, and Sparse contains tables with sparse content. These variations make it challenging to accurately predict the HTML table structure tags and bbox surrounding the cell content. Without SSP, the model performance suffers, and increasing the model complexity from base to large barely improves the performance. Instead, after pretraining the visual encoder on 1M tabular images, both the base and the large models have an average increase of 14.40 pp. Note the performance also scales along with the pretraining dataset size. When the pretraining dataset increases from 1M to 2M, the average performance continues to increase by 0.74 pp. Finally, all large models have an average performance gain of 1.51 pp over the base models.
@@ 5.2 Generalization of the Unified Training Objective
Table 3: UniTable’s unified training objective applies to both linear projection Transformer and hybrid CNN-Transformer architectures conventionally used in TR. Results on all four subsets of the SynthTabNet for table structure prediction evaluated with S-TEDS, and the same conclusion also applies to other tasks as elaborated in Appendix D.
@@
We showcase that our unified training objective, language modeling, not only works on models with linear projection Transformer, but also works with the hybrid CNN-Transformer architecture used in the TR literature. Table 3 presents the results on all four subsets of the SynthTabNet similar to the settings in Sec. 5.1. The main difference is that the linear projection in the visual encoder is replaced by a ResNet-18 CNN backbone. Such a modification leads to an increase of 12M parameters. The performance of this hybrid CNN-Transformer architecture is roughly on par with the linear projection model initialized from the SSP 1M tabular images, which shows that our training objective is agnostic to the choice of architecture. The significant performance gap between hybrid CNN-Transformer and linear projection Transformer trained from scratch verifies the observation identified in the previous literature [30]. Though hybrid CNN-Transformer has also achieved competitive results, we still recommend using the linear projection Transformer because of 1) capability of leveraging the power of SSP, 2) architectural compliance with VLM in natural image domain, and 3) the performance of hybrid CNN-Transformer is still worse than the SSP on 2M tabular images even with more total parameters. Though prior work has started to propose SSP for hybrid CNN-Transformer, this direction is still in its early stage as it is challenging in heavy computational cost and pretraining-finetuning discrepancy [9]. Moreover, most work only demonstrates the capability in discriminative tasks rather than the generative tasks used in the TR domain. Besides, existing SOTA VLMs, e.g., BLIP-2 [17] and LLaVA [20], still employ the linear projection and leave the Transformer to learn the interactions between different patches. As we believe the compliance of architecture is an important step towards generic VLMs, our success in replacing the CNN backbone with linear projection and leveraging a task-agnostic language modeling loss is a cornerstone of incorporating TR in modern VLM training.
@@ 5.3 Why does SSP work?
((a)) Background patches (red)
((b)) Patches for separations
@@
Figure 3: The key reason that SSP works is because each tokens have visual semantics and the codebook shows a fine-grained categorization to represent the implicit conventions in the table. The codebook used in SSP has learned to represent abstract concepts by using different groups of tokens to represent different concepts: (a) empty background and (b) separations within a table. Red highlights the token indices under investigation. Appendix E provides a zoomed-in version of these images labeled with token indices. We lay the color patches over the original tabular image to present the selected token indices from the 2M VQ-VAE.
@@
During SSP, the visual encoder is trained to fill in the masked tabular image by selecting visual tokens from the VQ-VAE codebook. Thus, we visualize the selected token indices and analyze whether the tokens have visual semantics. Fig. 3 presents the original table overlaid by the selected token indices from the 2M VQ-VAE codebook. We highlight the token indices under investigation in red. Appendix E provides a zoomed-in version of these images labeled with token indices. The input image size is $448\times 448$ and the patch size is $16\times 16$, so each image has $28\times 28$ tokens chosen from the 16384 entries in the codebook. First, we highlight the tokens representing the blank background (Fig 3a), which covers most of the empty space in the table; leaving only texts and separation lines nonhighlited. Then, we highlight the tokens representing the separation lines (Fig 3b), showing that they well-separate the header row, and the shaded and non-shaded rows. Taking a closer look at the token indices in Fig. 5 and 6 in Appendix E, we find the codebook has learned to represent an abstract concept by assigning multiple tokens, where each reflects a different scenario. For example, token “111” represents the separation above the gray color shading and token “964” represents the one below, and token “15282” represents the separation above a bold horizontal line and “10807” represents the one below. A certain type of separation also has multiple choices depending on the portion of the line within the patch, e.g., token “14181”, “8714”, and “10807” all representing the “below a bold horizontal line”, but the differences lie in the line thickness and amount of black inside the image patch. Such a fine-grained categorization also explains why the codebook needs so many tokens to represent the implicit conventions created by humans in the table.